Suppose the consumption equation is represented by the following :c= 250+0.75yd.

Calculate private savings


1
Expert's answer
2022-03-24T16:14:08-0400

Solution

c= 250+0.75yd.

Therefore,

Private Savings = National Income – Tax – Consumption


If saving =0, then C=Y

Y=250+0.75Yd

Y = 250+0.75(Y -40-0.2Y)

Y=250+0.75Y-30-0.15Y

Put like terms together and simplify

Y-0.6Y=220

0.4Y=220

"Y=\frac{220}{0.4}"

"Y=550"


Savings will be 550